
For some reason, the RA offset correction did not work for about ten minutes when the observing support tried to apply a correction.
At 4:27 UT, when we asked for Master Control remotelly from La Serena, The Interlocl and OCS lights became red. Then, when Alberto gave us control, they became green and the exposure that was aborted was restarted.
At 04:33 UT, the next images in the queue didn't require a slew, but SISPI too very long to start exposing.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2019-12-13 | 7.25 | 0.0 | 0.0 | 0.0 | 0.0 | 7.25 |
Links
[1] mailto:shenyue@illinois.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/28083/delete?destination=print/content/enr-ctio-blanco-4-m-2019-12-13
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/28083/edit?destination=print/content/enr-ctio-blanco-4-m-2019-12-13
[4] mailto:alfredo@noao.ctio.edu
[5] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/28084/delete?destination=print/content/enr-ctio-blanco-4-m-2019-12-13
[6] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/28084/edit?destination=print/content/enr-ctio-blanco-4-m-2019-12-13
[7] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/16730?destination=print/content/enr-ctio-blanco-4-m-2019-12-13
None.